A radio set is retuned from the 108MHz to 90.5MHz station using the variable capacitor.By what factor is the capacitance changed


1
Expert's answer
2021-02-07T19:08:17-0500

Let "f_1 = 108\times 10^6Hz", "f_2 = 90.5\times10^6Hz". The formulas for resonant frequencies are:


"f_1 = \dfrac{1}{2\pi\sqrt{C_1L}}\\
f_2 = \dfrac{1}{2\pi\sqrt{C_2L}}"

where "C_1, C_2" are corresponding capacitances. Dividing the first expression by the second and expressing ratio "C_2\/C_1", obtain:


"\dfrac{C_2}{C_1} = \left( \dfrac{f_1}{f_2} \right)^2\\
\dfrac{C_2}{C_1} = \left( \dfrac{108\times 10^6}{90.5\times10^6} \right)^2 \approx 1.42"

Answer. The capacitance changed by a factor of 1.42.